OpenStudy (anonymous):

Elaine is 7 years older than Jane. Their combines age is 23. How old is Elaine?

OpenStudy (alexandervonhumboldt2):

elanine=kane+7 elani+jane=23

OpenStudy (alexandervonhumboldt2):

substitue jane+7 or elanie: (jane+7)+jane=23 2jane=16 jane=8

OpenStudy (alexandervonhumboldt2):

jane=8

OpenStudy (alexandervonhumboldt2):

elanie=8+7=???????

OpenStudy (anonymous):

Thank you for explaining it! ^-^